OpenStudy (anonymous):

solve equation find exact solution if possible a)54^-x=1/25 b)e^(3x/4)=10 c)3^(2x)-3x-6=0

OpenStudy (anonymous):

for a I got 2.3 and for b I got 3/4 are they right? and I need help with c...

OpenStudy (campbell_st):

so is the 1st question \[54^{-x} = \frac{1}{25}\]

OpenStudy (anonymous):

ya

OpenStudy (campbell_st):

well I know for sure and certain is not 2.3

OpenStudy (campbell_st):

so do you know about logs...?

OpenStudy (anonymous):

ya

OpenStudy (campbell_st):

ok... so take the log of both sides and you get \[\ln(54^{-x}) = \log(\frac{1}{25})\] applying the log laws for powers you get \[-x \ln(54) = \ln(\frac{1}{25})\] now solve for x... what do you get..?

OpenStudy (anonymous):

so ln of each side ?

OpenStudy (campbell_st):

yep... that the 1st step then apply the log law for powers \[\log_{a}(b^x) = x \log_{a}(b)\]

OpenStudy (anonymous):

3.2 ?

OpenStudy (campbell_st):

nope... next step... divide both sides by ln(54) \[-x = \frac{\ln(\frac{1}{25})}{\ln(54)}\]

OpenStudy (anonymous):

so I make itt to multipl?

OpenStudy (campbell_st):

nope just do the division shown...

OpenStudy (campbell_st):

last step is to multiply both sides of the equation by -1 to find the value of x

OpenStudy (anonymous):

so doing div I get -.806

OpenStudy (campbell_st):

so did you multiply by -1 to get the value of x..?

OpenStudy (campbell_st):

I had -x = -0.806

OpenStudy (campbell_st):

so I feel x = 0.806 to check find the value of \[54^{-0.806} \] does it equal 1/25 or 0.04...?

OpenStudy (anonymous):

so the final answer is + 0.806

OpenStudy (campbell_st):

yep... that's my best guess...

OpenStudy (campbell_st):

want to take about b..?

OpenStudy (anonymous):

yes plz b and c

OpenStudy (anonymous):

for b I did Ln10=ln^3x/4 3x/4 = 4/3

OpenStudy (campbell_st):

ok... when you have the same base... and take the log, you just get the power \[\log_{a}(a^x) = x\] so take the base e log or ln of both sides and you get \[\frac{3x}{4} = \ln(10)\] so to find x, multiply both sides of the equation by 4 then divide both sides by 3

OpenStudy (anonymous):

so if I multiply both by 4 its 3x=ln40 ?

OpenStudy (anonymous):

2.5 ?

OpenStudy (anonymous):

ok so i got 1.22 @campbell_st

OpenStudy (campbell_st):

nope if you multiply by 4 you get \[3x = 4 \times \ln(10)\] then divide by 3

OpenStudy (anonymous):

so x=4/3

OpenStudy (campbell_st):

yep \[x = 4 \times \ln(10) \div 3\]

OpenStudy (anonymous):

ok :D so how about c do I add the 6 to both sides first?

OpenStudy (campbell_st):

they want an exact solution for all questions so a) \[x =- \ln(\frac{1}{25}) \div \ln(54) = -\frac{\ln(1) - \ln(25)}{\ln(54)}\] and b) \[x = \frac{4\ln(10)}{3}\]

OpenStudy (anonymous):

ok/

OpenStudy (campbell_st):

is the last equation \[3^{2x} -3^{x} - 6 = 0\]

OpenStudy (anonymous):

yup

OpenStudy (campbell_st):

ok... this is basically a quadratic equation that can be factored if I let \[u = 3^x\] the equation becomes \[u^2 - u - 6 = 0\] which can be factored to \[(u - 3)(u + 2) = 0\] so then the solutions are u = 3 and u = -2 doing the reverse substitution \[3^x = -2\] has no solution... so the only equation you need to solve is \[3^x = 3^1\] what do you think x is...?

OpenStudy (anonymous):

1?

OpenStudy (campbell_st):

thats it... all done

OpenStudy (anonymous):

:D thank you

OpenStudy (campbell_st):

glad to help